UARS Weekly Status Report

19 January 2001




********************************************************************

                UARS WEEKLY STATUS REPORT - 19 JAN 2001

********************************************************************


	>>>>> An Update on UARS Attitude Control <<<<<

Attitude of the UARS spacecraft is being controlled by 3-axis
magnetometer (TAM) data which are being processed by the on-board
computer.  The plan is to rely on on-board processing of TAM data
for attitude control with the Fine Sun Sensor (FSS) in monitor
mode.  However, at the high solar beta angles which are being
encountered between 16 January and 25 January the Sun is out of
the Fine Sun Sensor field of view.  Our indications are that S/C
attitude is being maintained within specifications using only the
3-axis magnetometer.  The FOT would welcome any feedback from
instrument teams about spacecraft attitude as they have observed
it.

***********
SOLSTICE
***********

SOLSTICE is in good health, collecting solar data on all available orbital
tracking periods and stellar calibration observations whenever possible. We
continue to gather enough data to meet all the SOLSTICE science objectives
during this period of reduced tape recorder operations.

***********
SUSIM
***********

SUSIM performed its daily solar scans of the sun during periods of
predicted UARS telemetry.

***********
ACRIM
***********

Nothing received.

***********
PEM
***********

Nothing received.

***********
HALOE 
***********

Nothing received.

***********
MLS
***********

Nothing received.

***********
HRDI
***********

HRDI continues to operate normally. 

***********
WINDII
***********

Nothing received.

***********
MPG
***********

The instrument operations schedule for the next yaw period, 
9 February to 15 March, has been distributed.  Daily 
processing of instrument Daily Activity Plans continues
with no problems. 

***********
CDHF
***********

Nothing received.
                  
***********
FOT
***********


                         REPORTING PERIOD

This report covers 06 January 2001 (Orbit 50955, GMT Day 006) through 12 
January 2001 (Orbit 51060, GMT Day 012).

                       SPACECRAFT OPERATION

The observatory is now in Forward Flight (Southern Hemisphere viewing) and is 
performing nominally under two-battery operations.  The instrument operational 
changes were:


                  INSTRUMENT OPERATIONAL CHANGES


 Instrument     Time       Orbit     Comment
 PEM          011/0014     51031     AXIS ON

The beta angle ranged between the angles of 30.4 degrees to 51.7 degrees for 
this report period.  The beta angle is now increasing towards a maximum angle 
of 77 degrees on 21 January 2001 (DOY 021).

Spacecraft battery 1, 2, and 3 performance monitoring continues.  Battery 1 
remains OFF the charge relay.  Battery 2 and Battery 3 are currently 
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.  Battery 
temperatures are stable with a temperature delta between Battery 2 and Battery 
3 of 2.2 and 1.9 degrees C respectively.  Additional battery performance data 
for each battery on 05 January 2001 and 12 January 2001 are provided in 
Appendix D.

The clock error ranged from 12 to 12.4 msec during this report period. 
Clock rate adjustments are listed in Appendix C.

All instrument and subsystem engineers were advised on the initial excursion of 
ALERTS or Out-Of-Limits conditions detected by the FOT for this report.  A 
summary is listed below and the details of ALERT and Out-Of-Limits occurrences 
are listed in Appendix B.

     OUT-OF-LIMITS  =  01
     ALERTS         =  00

The following real-time command sheets were executed for the listed operational 
element(s) during this report period (see Appendix A).

     FOT           =  27
     PEM           =  01


UARS Anomaly Reports initiated during this report period are listed in Appendix 
E and summarized below.

     FOT           =  01

Four (4) Anomaly Report remained OPEN at the end of this report period (see 
Appendix E).

TRANSPONDER FREQUENCIES:
 Transponder A:  00/315  -  2287.496493
 Transponder B:  00/314  -  2287.498584


                     GROUND SYSTEM OPERATION

One (1) operational support problems occurred this period.  Anomaly Reports are 
included in Appendix E.

The data loss calculations supplied by Data Capture Facility (DCF) from mission 
start through UTC Day 01/011 show the observatory data loss to be 300 hours, 08 
minutes, 41.256 seconds. This shows an overall increase of 0 seconds since the 
last report period This is a 0.39 percent data loss, which equals a 99.61 
percent data capture for the mission.

 There was no data loss for the week.

                   OTHER SIGNIFICANT EVENTS

                   UARS SATELLITE OPERATIONS

UARS continues to maintain attitude control on TAM alone this week.  The FSW 
patch of the nonorthogonal alignment was successfully uplinked on 10 January 
2001. The patch improved the TAM's performance, as expected, and UARS is set 
to fly through the high beta period with the TAM as its only input to UFLTR.  
More fine tuning of the TAM is needed before it will be operating at its 
maximum effectiveness, but it has proven to be stable enough to maintain all 
science requirements.


                            MEETINGS

The FOT attended a TAM Attitude Meeting on 08 January 2001.
The FOT attended a TAM Software Patch Meeting on 09 January 2001.
The FOT attended a TAM Attitude Meeting on 10 January 2001.
The FOT attended a UARS Retrieval Meeting on 10 January 2001.


                    FUTURE SIGNIFICANT EVENTS


The next UARS Yaw slew is scheduled for 09 February 2001 (DOY 040).  This 
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere 
viewing).


Attachments: Appendix A,  Real-time Commands Sheets
             Appendix B,  Out-Of-Limits & ALERT Occurrences
             Appendix C,  Clock & Solar Array Rate Corrections
             Appendix D,  UARS Battery Performance Data
             Appendix E,  UARS Anomaly Reports

                           APPENDIX A
                    REAL-TIME COMMAND SHEETS
                   Week ending 12 January 2001

ORBIT     AOS         NO.      SUBSYS     REASON
-----     ---         ---      ------     ------
50958     006/0409    01-021   OBC        Clk Adj 0.0 to -20.6 msec/day
50960     006/0708    01-020   NBTR       NOOPS D006
50960     006/0738    01-022   OBC        Clk Adj -20.6 to 20.6 msec/day
50965     006/1611    01-023   ACS        TAM ALIGNMENT
50967     006/1800    01-024   NBTR       NOOPS D007
50979     007/1348    01-025   OBC        Clk Adj 20.6 to 0.0 msec/day
50982     007/1808    01-026   NBTR       NOOPS D008
50990     008/0711    01-026   NBTR       NOOPS D008
50994     008/1433    01-027   ACS        TAM ALIGNMENT
50997     008/1909    01-028   NBTR       NOOPS D009
51000     009/0011    01-029   ACS        MAGMOD Time Correction
51005     009/0712    01-028   NBTR       NOOPS D009
51009     009/1352    01-030   ACS        TAM ALIGNMENT
51009     009/1434    01-031   POCC       FILE XFER
51012     009/1910    01-032   ACS        TAM ALIGNMENT
51012     009/1910    01-033   NBTR       NOOPS D010
51020     010/0750    01-033   NBTR       NOOPS D010
51024     010/1354    01-034   FSW        FSW Update for TAM
51025     010/1544    01-035   OBC        Reset Status Buffer
51027     010/1911    01-036   NBTR       NOOPS D011
51031     011/0014    01-037   PEM        AXIS POWER ON
51032     011/0215    01-038   OBC        Clk Adj 0.0 to 20.6 msec/day
51035     011/0722    01-036   NBTR       NOOPS D011
51041     011/1731    01-039   OBC        TAM ALIGNMENT
51042     011/1911    01-040   NBTR       NOOPS D012
51050     012/0716    01-040   NBTR       NOOPS D012
51055     012/1439    01-041   ACS        TAM ALIGNMENT
51057     012/1913    01-042   NBTR       NOOPS D013
__________________________________________________________________________

                           APPENDIX B
             OUT-OF-LIMITS & ALERTS OCCURRENCES
                Week ending 12 January 2001

The following Out-Of-Limits occurrences were detected:

MNEMONIC     STATE     DESCRIPTION                  ORBITS
--------     -----     -----------                  ------
HATBAT       Y-HI      HALOE Bolometer Area Temp    51056-60


The following ALERT occurrences were detected:

MNEMONIC     DESCRIPTION     ORBITS
--------     -----------     ------
-None this report period
_______________________________________________________________________________

                           APPENDIX C
            CLOCK & SOLAR ARRAY RATE CORRECTIONS
                Week ending 12 January 2001

                  CLOCK DRIFT RATE ADJUSTMENTS

ORBIT     TIME          CLOCK RATE CHANGE
-----     -----         -----------------
50958     006/0410      Clk Adj 0.0 to -20.6 msec/day
50960     006/0740      Clk Adj -20.6 to 20.6 msec/day
50979     007/1354      Clk Adj 20.6 to 0.0 msec/day
51032     011/0210      Clk Adj 0.0 to 20.6 msec/day

___________________________________________________________________________

                           APPENDIX D
                UARS BATTERY PERFORMANCE DATA
                  Week ending 12 January 2001


     05 Jan 2001 - GMT Day 005 Beta = 30.4 deg, SA Pos = 269 deg

     DIFF V (mV)    TEMP (C)     CURR (amp)   EON V   AVE C/D   AVE
BAT  MAX/MIN        MAX/MIN      MAX/MIN      MIN     RATIO     DOD(%)
---  -----------    ----------   -----------  ------  -------   ------
1    +728.0/+728.0  -3.05/-4.31  +0.0/+0.0    14.2    0.000     0.0
2    +0.0/-44.8     +5.47/+4.51  +16.0/-15.2  25.9    1.030     21.0
3    +11.2/-22.4    +4.51/+3.55  +15.6/-15.2  25.8    1.030     20.9


     12 Jan 2001 - GMT Day 012 Beta = 51.7 deg, SA Pos = 269 deg

     DIFF V (mV)    TEMP (C)     CURR (amp)   EON V   AVE C/D   AVE
BAT  MAX/MIN        MAX/MIN      MAX/MIN      MIN     RATIO     DOD(%)
---  ------------   ----------   ----------   ------  -------   ------
1    +728.0/+728.0  -3.68/-4.94  +0.0/+0.0    14.2    0.000     0.0
2    +11.2/-95.2    +5.15/+2.92  +25.6/-15.2  25.1    1.050     19.9
3    +22.4/-44.8    +3.87/+1.97  +25.6/-15.6  25.1    1.040     19.7


     V/T LEVEL CHANGES
ORB#           TIME          LVL FR-TO
-----          ----------     ----------
-no changes    V/T 5

__________________________________________________________________________
                              
                                APPENDIX E
                   UARS ANOMALY REPORTS
                 Week ending 12 January 2001

                  NEW ANOMALY REPORTS GENERATED
AR NO.     SUBSYS    ORBIT     TIME      TTR      PROBLEM/DATA LOSS
------     ------    -----     ----      ---      -----------------
01-002     FOT       51005     009/0749  14170    Generic Late Acquisition

            ANOMALY CLOSURE INFORMATION RECEIVED
AR NO.     SUBSYS     ORBIT     TIME     TTR      PROBLEM/DATA LOSS
------     ------     -----     ----     ---      -----------------


           ANOMALY INVESTIGATION REPORTS GENERATED
AR NO.     SUBSYS     ORBIT     TIME     TTR      PROBLEM/DATA LOSS
------     ------     -----     ----     ---      -----------------
     -None this report period


                       OPEN ANOMALY REPORTS
AR NO.     SUBSYS     ORBIT     TIME      TTR      PROBLEM/DATA LOSS
------     ------     -----     ----      ---      -----------------
00-053     FOT        50168     319/1113  23157    Late Acquisition
00-054     FOT        50280     326/2147  23187    RESET PTP /2:39 32K
00-056     FOT        50584     347/0502  23302    Premature LOS
01-002     FOT        see list  Jan 01    14170    Generic Late Acquisition

	 	SEVNTFW01/02